CRYSTALC:

Riddle- in two years i will be twice as old as I was 5 years ago how old am i

snowflake0531:

12

carmelle:

12; 5 x 2 + 2

CRYSTALC:

thanks

BigDawhg:

12

jhonyy9:

what may be the equation to solve this ?

SmokeyBrown:

Current age = x x + 2 = age in 2 years x - 5 = age 5 years ago x + 2 = 2(x-5) ; (age in 2 years is twice age 5 years ago) x + 2 = 2x - 10 12 = x

Extrinix:

@jhonyy9 wrote:
what may be the equation to solve this ?
It's a riddle, amicus meus.

jhonyy9:

ok i know but to solve it you need write an equation

sxdsouls:

\[5~x ~2 + 2 = 12\]